# From Infants to AI: Incorporating Infant-like Learning in Models Boosts Efficiency and Generalization in Learning Social Prediction Tasks

### v2

@@ Results
Experiment 4: Goal Attribution
In this experiment, a related yet distinct concept, ’Goal,’ is investigated. Similar to the idea of actor type, it is assumed, based on the empirical literature, that infants can attribute goals to animate entities, based on cues such as certain types of contact or gaze direction (?, ?, ?). Similar to the ’animacy’ concept, the concept of ’having a goal’ is provided to the cognitive model, and the results are compared with a naive model that can infer the goal, but is not directly provided with the concept.
Data
@@
For the cognitive model, the goal is provided as part of the actor’s representation (e.g.: ’girl with goal cake’), but its implications are not specified and must be learned during training. In contrast, the naive model must infer the goal directly from the data. The only difference between the models is the inclusion of the goal representation in the cognitive model. An example of the goal attribution representation and training data can be found in the Appendix.
@@
A Note on Vision-and-Language Models
Recent Vision-and-Language Models (VLMs) combine two modalities, mirroring, to some extent, how humans receive input from multiple sources and senses. These models can be considered ’hybrid’ in relation to our work—naive yet trained on vast datasets that resemble cognitive behavior. The extent to which these models genuinely mimic human behavior and integrate modalities to enhance human-like abilities remains an active area of study.
Although VLMs are not central to our research reported here, we briefly investigated their performance. We prompted a VLM (Gemini (?, ?)) with 208 three-frame image sets and posed various questions to assess whether its behavior could be considered ’as smart as infants’ in a Woodward-like task. Our findings reveal that the VLM showed no significant deviations from random selection, performing at chance levels on this task. The VLM did not show any preferred behaviour for either animate actors (following the same goal) or non-animate actors (following the same trajectory). This underscores that, despite significant progress, these models lack key aspects of understanding that infants possess. Furthermore, the explanations provided by the model failed to demonstrate the depth of understanding characteristic of infant cognition. Additional details of this investigation can be found in the Appendix.
Discussion
The current study examined a basic characteristic of early human learning, by modeling some aspects of how early-acquired concepts are used in the learning of subsequent concepts and comparing the results with standard deep network modeling. The results show that the use of early concepts in the learning of new concepts can lead to higher accuracy and more efficient learning. The results of comparing generalization to new tasks also suggest that the integration of early and new concepts during the learning process can shape the representation of the concepts acquired by the model. Although the dataset is simplified compared to natural images, prior work has shown that infants succeed in tasks with similarly simplified stimuli (?, ?).
@@ Discussion
What can cause the ‘cognitive’ models to learn better than the ‘naïve’ ones? There is empirical and theoretical evidence suggesting that the difficulty of learning a function F by a deep network model depends on the complexity of F: the learning will require more data, and the probability of finding F, or a close approximation, starting from a random initialization, will decrease (?, ?). If this notion is correct, then learning by the cognitive network, where the learning process is divided into components, may help the learning by decomposing the overall task into components, where the complexity of each component is lower than the complexity of the full task. The general suggestion raised by this possibility is that the training of models could benefit from some form of hierarchical decomposition, where simple concepts are learned first and can be used in the learning of more complex concepts.
Current large models are trained on huge data sets, and they strive to create so-called foundational models, which can deal with a broad range of tasks. Such models have achieved impressive results. However, multiple studies have shown examples where the learning by current large models appears to be different from human leaning; they sometimes ‘show failures on surprisingly trivial problems’ (?, ?), or ‘show basic errors in understanding that would not be expected even in non-expert humans’ (?, ?). It is possible that shortcomings of this kind will be overcome by additional data and training. Alternatively, it may prove useful to combine current large models and human learning in some useful manner. For example, it may be possible to learn early concepts by adapting methods similar to self-supervised learning of early concepts in humans, and use early-learned concepts in the learning of more complex ones. If such training proves feasible, it would be of interest to test the effects of making the model more human-like both on its performance, as well as its similarity to human behavior in different visual tasks.
Acknowledgments
ST acknowledges support from the Ariane de Rothschild Women Doctoral Program. SU acknowledges the Weizmann Institute of Science for its support.
Appendix A Appendix
Experiments
Technical Details
All experiments were conducted on up to four NVIDIA RTX-6000 GPUs, each with 24GB of RAM. In the Scene Representation part, we employed the AdamW optimizer (?, ?) with a learning rate of $2\times 10^{-5}$ and a weight decay of 0.01, using a batch size of 10. For the Prediction Generation part, we used the Adam optimizer (?, ?) with a learning rate of $10^{-6}$, with other parameters determined by the model. Training was conducted with a batch size of 40. Hyperparameters were selected without an exhaustive tuning process. We did not explicitly set a random seed, allowing the environment to generate one for each run.
Experiment 4: Goal Attribution
@@ Appendix A Appendix
Figure 9 illustrates the representation and data used in the goal attribution experiment.
@@
Figure 9: Example representation (left) and training data (right) from the goal attribution experiment. The cognitive concept is highlighted in bold and purple. The right image shows an example of training data, where each actor with a goal is presented in 10 repetitions with varying distractors across examples.
A Note on Vision-and-Language Models
Data
An example of the VLM data is shown in Figure 10. We generated four sequences for each triplet of an actor and objects to ensure the data was counterbalanced. Two types of questions were used to balance the data in terms of animate and non-animate entities (see Figure 10).
The first question was: ”This is a sequence of images representing a scene from real life. What do you predict the ’ACTOR’ will grasp in the next unseen frame?” The second question was: ”This is a sequence of images representing a scene from real life. Where do you predict the ’ACTOR’ will be in the next unseen frame?”
We used these two versions because the words ”what” and ”grasp” may be more associated with animate beings, while ”where” and ”will be” may be more linked to locations for non-animate entities. This choice helped us avoid potential bias.
A total of eight sequence-question pairs were created for each actor, with 11 animate and 11 non-animate actors, resulting in 176 sets. We probed Gemini (?, ?) with each sequence and analyzed the results.
Figure 10: VLM Data. The left figure shows an example of the four sequences generated for each triplet of an actor and two objects, ensuring counterbalanced data. The right figure presents the two types of questions: the first being more associated with animate behavior and the second with non-animate behavior.
Analysis
We first analyzed the data by considering all correct answers (i.e., cases where Gemini’s response correctly identified the expected move by the actor) and labeled the rest as incorrect. A t-test was conducted to compare these results to a Bernoulli distribution, representing random selection. The results, shown in Table 1, indicated significantly more incorrect responses than correct ones.
Next, we examined sub-categories such as agent type (animate or non-animate) and question type. We found no significant differences between correct and incorrect responses in most sub-categories, except for Question 2, where incorrect responses were significantly more frequent. We hypothesized that this could be due to irrelevant or hallucinated answers (e.g., irrelevant location between the two objects or unrecognized objects leading to fabricated responses).
To focus specifically on the difference between correct and incorrect responses, we removed irrelevant and hallucinated responses, leaving 147 results. Another t-test was conducted, as shown in Table 2, which revealed no significant difference between the two groups. Subsequent analysis of the sub-categories also showed no significant differences. We concluded that Gemini performed at about chance level on these tasks.
Fraction p-value
Correct Responses 0.417 0.023
Incorrect Responses 0.585
Table 1: VLM First Analysis. This table presents the analysis of all correct and incorrect responses from Gemini. The fractions represent the proportion of each category (correct/incorrect) relative to the total number of responses. The p-value indicates the significance of the t-test.
Fraction p-value
Correct Responses 0.496 0.934
Incorrect Responses 0.503
Table 2: VLM Second Analysis. This table presents the analysis of correct and incorrect responses from Gemini after excluding irrelevant and hallucinated responses. The fractions represent the proportion of each category (correct/incorrect) relative to the total number of responses. The p-value indicates the significance of the t-test.

### v3

@@ 4.3.2 Results
4.4 Experiment 4: Goal Attribution
In this experiment, a related yet distinct concept, ’Goal,’ is investigated. Similar to the idea of actor type, it is assumed, based on the empirical literature, that infants can attribute goals to animate entities, based on cues such as certain types of contact or gaze direction [42, 30]. Similar to the ’animacy’ concept, the concept of ’having a goal’ is provided to the cognitive model, and the results are compared with a naive model that can infer the goal, but is not directly provided with the concept.
4.4.1 Data
@@
For the cognitive model, the goal is provided as part of the actor’s representation (e.g.: ’girl with goal cake’), but its implications are not specified and must be learned during training. In contrast, the naive model must infer the goal directly from the data. The only difference between the models is the inclusion of the goal representation in the cognitive model. An example of the goal attribution representation and training data can be found in Appendix A.2.
@@ 4.4.2 Results
5 Do Large Vision–Language Models Exhibit Human-Like Behavior in the Woodward Task?
Learning in humans and in our cognitive model begins with simple concepts that guide prediction and shape the final representations. In contrast, current Vision–Language models (VLMs) and Multi-Modal Large Language Models (MMLLMs) are typically trained end-to-end with the goal of optimizing performance in a final task. In this section, we examine whether VLMs acquire the core concepts that are part of the human cognitive system as part of their training.
Specifically, we examined whether the behavior of VLMs resembles human behavior in the Woodward task. In humans, predictions in the Woodward configuration are consistent with the notion that animate agents move toward their preferred object targets, whereas inanimate objects do not show such a preference, but show some preference to repeat their previous trajectories. We tested whether VLMs exhibit similar predictive behavior, and our evaluation focused on the performance of VLMs following their general training, without specific fine-tuning for the Woodward task.
Testing VLMs on the Woodward-type tasks is a part of a broader comparison, since object-tracking and trajectory-tracking behaviors, as studied in the Woodward task, are instances of a more general and fundamental distinction between the behavior of animate versus inanimate entities. The behavior of animate entities is determined by their internal goals, preferences, and beliefs, whereas non-animate behavior is determined by external forces and circumstances. For humans, this is an intuitive and basic dichotomy between ‘intuitive physics’ and ‘intuitive psychology’ [21], which begins to emerge already at early development [36]. We describe below our comparison of human subjects and VLM models in a version of the Woodward task. As described in detail below, we compare their capacity to correctly predict the behavior of animate and non-animate agents in the Woodward test scenario, with the main focus on predicting that animate agents prefer object-tracking (preference for moving towards the preferred object) over trajectory-tracking (preference to repeat a previous trajectory).
5.1 Human testing
Figure 11: Example data from the human experiment and VLM testing. Top: Animate condition. Bottom: Non-animate condition. Left: Example frames from the videos presented to human participants and models. Right: An example of the question presented after the video.
5.1.1 Participants
We recruited 200 participants through the Prolific research platform. All participants provided their informed consent and were compensated for their participation. Participants completed the experiment online.
5.1.2 Method
To test human behavior in Woodward-like tasks, we designed two types of video conditions: animate and non-animate (see Figure 11). The dataset included 100 animate videos and 100 inanimate videos. Each participant viewed a single video.
In the animate condition, each video begins with two goal objects located at the top of the screen. An animate actor (e.g., a girl or a dog) then enters from the bottom of the screen and moves toward one of the objects until reaching it. A gray screen is then shown, after which the two objects reappear with their locations swapped. After the video ends, participants are shown two images and asked to predict which one is more likely to be the next part of the video. In one image, the actor reaches toward the same goal object, which is now in a new location. In the other image, the actor reaches toward the same location as before, which now contains the other object.
In the non-animate condition, the videos show two objects at the top of the screen and a non-animate actor (e.g., a ball) at the bottom. Next to the non-animate actor there is a pendulum, which provides the physical cause of the motion by pushing the actor and causing it to move in one direction on the screen and stop near one of the objects. A gray screen is then shown, and the two objects reappear with their locations swapped. The prediction task is similar to the animate condition. The participants see two images, showing the actor at the bottom near the pendulum and the two objects at the top. In each image, the non-animate actor is positioned near one of the sides, and participants must choose which image is more likely to represent the continuation of the video. Participants were also asked to provide an explanation for their prediction after choosing one of the images. See Appendix B for additional details about the human experiment.
5.1.3 Results
Out of 100 participants in the animate condition, 75 answered correctly, and 81 out of 100 participants answered correctly in the non-animate condition. Unlike the original Woodward experiment, which was conducted with infants, our experiment was conducted with adults. The results in the animate condition are consistent with the findings of the original study. In contrast, in the original Woodward experiments the non-animate condition was close to chance level, and this preference develops only later.
5.1.4 VLM Testing
Humans Gemini 3.1 Flash Gemini 3.1 Pro Claude 3.5 Sonnet Claude Opus 4.6
Animate 75 41 63 53 37
Non-animate 81 47 90 50 94
[…]
Humans GPT-4o GPT-5.2 Qwen VL Max Qwen 3.5 Plus
Animate 75 35 84 48 85
Non-animate 81 60 58 49 88
[…]
Table 1: Accuracy of humans and vision-language models on the animate and non-animate conditions (N = 100 trials per condition). Models marked with were significantly different from human performance (two-proportion z-test, $\alpha=0.05$).
5.1.5 Models
To evaluate model behavior in Woodward-like tasks, we benchmarked eight models: two Gemini models, two GPT models, two Qwen models, and two Claude models. For each model family, we selected two versions: the most recent available model and an earlier-generation version.
5.1.6 Method
The VLM test was similar to the human experiment and used the same data. The models were presented with the videos and prompted to predict the next step by choosing between two possible images (A or B). This procedure was applied to both the animate and non-animate conditions. The models were instructed to explicitly provide a prediction (A/B) along with an explanation. Additional details and statistical analysis details are provided in Appendix C
5.1.7 Results
We compared the models’ predictions with the ground truth for both conditions and across all models, see table 1. In general, the more recent model versions performed better than the earlier-generation models in both conditions. To assess alignment with human performance, we conducted two-proportion z-tests comparing each model to humans within each condition ($\alpha=0.05$). In the animate condition, most models performed significantly lower than humans. Gemini 3.1 Pro was lower than humans but did not differ significantly. GPT-5.2 and Qwen 3.5 Plus achieved higher accuracy than humans; however, these differences were not statistically significant. In the non-animate condition, most models again performed significantly lower than humans. Gemini 3.1 Pro and Qwen 3.5 Plus achieved higher accuracy than humans, but the differences were not statistically significant. Claude Opus 4.6 was the only model that performed significantly higher than humans in this condition.
The difference between the human system and the VLMs in the tasks described above is striking if we take into account the amount of data and training used in the learning process. The training of advanced LLMs requires a huge text data set, measured in up to trillions of text tokens (e.g. 15 trillions in Meta’s Llama) and billions of image-text pairs (e.g. 5.86 billions in LAION-5B). It is also notable that 6–9-month-old infants’ performance in the animate condition (approximately 75%; [42]) in a similar task is relatively high compared with the models we tested.
6 Discussion
The reported results of the current studies demonstrate significant effects of early cognitive concepts, leading to efficient learning and broad generalization. A natural question that arises is whether similar benefits would extend to more natural real-world applications, and to the training of current large models. In this section, we briefly discuss two fundamental issues related to the results and their implications. First, we discuss how early-emerging concepts, incorporated in a hierarchy of concepts, may contribute to better learning. Second, we consider the possibility of incorporating the equivalent of early human concepts into current vision and vision-language models.
6.1 Possible Contributions of the Early Cognitive Concepts to Better Learning
Two main factors appear to be the basis for the advantage of the cognitive models. The first has to do with the discovery and incorporation of useful and informative concepts. When we learn about a new domain, we often benefit from getting acquainted with concepts that capture basic entities and interactions in the domain; e.g., concepts such as supply, demand, equilibrium, incentives, and competition are useful for understanding aspects of the economy. The suggestion is that the discovery of useful concepts and their interactions can be a demanding part of the learning process; therefore, having them available can lead to more efficient learning and a better understanding of the domain. The concepts of ’Animacy’ and ’Goal’ discussed in the studies above were shown to be useful in capturing a basic distinction between two major categories of entities: animated agent vs. inanimate objects. The general conjecture is that the conceptual structure that emerges in early human visual learning, selected by evolutionary processes, can be useful in the learning and understanding of relevant domains.
@@ 6.1 Possible Contributions of the Early Cognitive Concepts to Better Learning
A second benefit of early cognitive concepts for learning a given task may also come from their role in decomposing the learning of the task into simpler components. There is empirical and theoretical evidence suggesting that the difficulty of learning a function F by a deep network model depends on the complexity of F. As complexity increases, learning is likely to require more data, and the probability of identifying F, or a close approximation, starting from a random initialization, is likely to decrease. Empirical support for these relations was shown for several complexity measures, e.g. in [41]. If this notion is correct, the training of a model may benefit from a decomposition of the target function into components, where the complexity of each component is lower than the complexity of the full task. This suggests that the cognitive model used in our study and similar models could benefit from a hierarchical decomposition, in which simple early concepts are learned on their own first and then used to support the learning of more complex concepts.
6.2 Possible Implications for Current Vision and Vision-Language Models
Our study has demonstrated that incorporating early cognitive structures can enhance the learning process, making it more efficient and accurate, and can lead to broader generalization, thereby facilitating a deeper understanding of the learned task. An important open question is whether the benefits shown in a simple example could also apply to large vision and vision-language models, and to complex real-world tasks. Current large models have achieved impressive performance, but they are still far from perfect. For example, models have been shown to sometimes ‘show failures in surprisingly trivial problems’ [7] or ‘show basic errors in understanding that would not be expected even in non-expert humans’ [28].
One common approach to improve existing models is to increase and improve the data used for training. A complementary approach may be based on the incorporation of useful early concepts within a hierarchical conceptual structure. Such concepts may be adopted from the human system or identified by computational methods. Our results showed examples in which the incorporation of early conceptual structure improved the model’s learning in terms of accuracy and the amount of data used for training. In addition to these improvements in the learning process, our results, together with previous experiments, suggest that in the learning of a new domain, the early human conceptual structure contributes to what can be considered as understanding of the new domain exhibited by the model.
An example of such an improved understanding comes from studies related to the Woodward task. Experiments on distinguishing animate and non-animate agents have shown that self-propulsion is a strong early cue for animate agency [31]. This property was used in a later study, in which infants at five months of age were shown a box that started to move across a table on its own (self-propulsion) [27]. Subsequently, they were shown a version of the Woodward test, where the moving agent was the box that they had seen previously. The test showed that the infants treated the box as an animated agent with the goal of reaching its preferred object. These results show that infants can identify novel objects as animate, and they link this attribute with the agent. The results further show that infants follow the implication of the agent’s attribute, leading them to expect the box to show a preference for object-following. In contrast, the large models we examined (Table 1), did not exhibit a similar association between identifying an entity as an animate agent and following the implication of this identification in terms of predicting the expected agent behavior. We tested the eight large models (in Table 1) on their ability to identify the agents we used in our studies as either animate or non-animate, and found that all the models reached high accuracy in this task (average identification accuracy of 98%). At the same time, for six of the eight models, the accuracy in correctly predicting the behavior of animate agents was low (between 35%-53%), indicating that, unlike infants, they did not follow the animacy implication.
In future research, it would be of interest to study the potential benefits of incorporating aspects of early human conceptual structures into network models. Many of the early conceptual structures acquired by infants are based on visual learning, and this early learning develops before substantial language acquisition. This suggests that attempts to incorporate early concepts into vision-language models could be directed, at least in part, toward the visual components of the models. This suggestion is also supported by previous studies, which have shown that the recognition of sophisticated relationships and interactions occurs in the visual system [14], and have found similarities between human and model processing of social interactions using videos similar to ours [35]. It will therefore be of interest to train the visual part of vision-language models on early visual tasks and basic concepts that are incorporated into the human system. Following training, it will be possible to test for possible improvements in the models. Improvements can fall into two categories: the first is improvement of the models’ capacity in different tasks, and the second is making the models more aligned with human cognition. If successful, the incorporation of early conceptual structures in visual processing could contribute to both the study of human vision and the improvement of future computer vision and multimodal models.
Declarations
Funding This work was supported by an Ariane de Rothschild Women Doctoral Funding awarded to Shify Treger. Additional support was provided by research funds from the Weizmann Institute to Shimon Ullman. Competing Interests The authors declare no competing interests. Ethics Approval The human participant study was approved by the Institutional Review Board (IRB) of the Weizmann Institute of Science, Rehovot, Israel. All procedures were carried out in accordance with the relevant guidelines and regulations. Informed Consent Informed consent was obtained from all participants prior to participation in the study. Consent for Publication A preliminary version of parts of this work appeared in the Proceedings of the Annual Meeting of the Cognitive Science Society (2025). Data Availability The datasets will be made publicly available upon acceptance of the article. Materials Availability Not applicable. Code Availability The code used for the experiments will be made publicly available upon acceptance of the article. Author Contributions S.T. and S.U. conceived and designed the research; S.T. conducted the experiments; S.T. and S.U. analyzed the data and wrote the paper; S.U. supervised the research.
Appendix A Experiments
A.1 Technical Details
All experiments were conducted on up to four NVIDIA RTX-6000 GPUs, each with 24GB of RAM.
In the Scene Representation stage, we employed the AdamW optimizer [26] with a learning rate of $2\times 10^{-5}$ and a weight decay of 0.01, using a batch size of 10.
For the Prediction Generation stage, we used the Adam optimizer [17] with a learning rate of $10^{-6}$; other optimization parameters were set to the model defaults. Training in this stage was conducted with a batch size of 40.
Hyperparameters were selected without an exhaustive tuning procedure. We did not explicitly fix a random seed, allowing the computational environment to generate a seed independently for each run.
A.2 Experiment 4: Goal Attribution
@@
Figure 12 illustrates the representation and training data used in the goal attribution experiment.
@@
Figure 12: Example representation (left) and training data (right) from the goal attribution experiment. The cognitive concept is highlighted in bold and purple. In the training data, each goal-directed actor appears in 10 repetitions with varying distractors across examples.
Appendix B Human Experiment
B.1 Participants
A total of $N=200$ participants were recruited via Prolific, with 100 participants assigned to each condition (animate and non-animate).
Each participant completed a single trial only, and no stimulus was repeated within a condition. Participants reported normal or corrected-to-normal vision, fluency in English, and completion of high school (or equivalent education).
B.2 Procedure
The experiment was conducted via Prolific. Participants were required to confirm normal or corrected-to-normal vision, completion of high school (or an equivalent educational program), fluency in English, and no prior participation in any of the experiments reported in this paper.
To ensure data quality, we included two attention checks: (1) participants were required to type a specific randomly generated number into a text box; (2) following the main task, participants were asked to name at least one object that appeared in the video.
Only two participants were excluded for failing to complete the task.
B.3 Stimuli for Human Experiment and VLM Testing
Stimuli consisted of icon-based visual scenes. Icons were obtained from the Flaticon website under appropriate licensing terms, and additional assets were generated using GPT- and Gemini-based image-generation interfaces.
Appendix C VLM Testing
C.1 Statistical Analysis Details
Gemini 3.1 Flash Gemini 3.1 Pro Claude 3.5 Sonnet Claude Opus 4.6
Animate $1.11\times 10^{-6}$ 0.0666 0.00119 $6.19\times 10^{-8}$
Non-Animate $5.48\times 10^{-7}$ 0.0707 $4.00\times 10^{-6}$ 0.00544
[…]
GPT-4o GPT-5.2 Qwen VL Max Qwen 3.5 Plus
Animate $1.31\times 10^{-8}$ 0.1149 $8.73\times 10^{-5}$ 0.0771
Non-animate 0.00113 $4.12\times 10^{-4}$ $2.10\times 10^{-6}$ 0.1714
[…]
Table 2: P-values from two-proportion z-tests comparing each model’s accuracy to human performance.
Human accuracy in each condition was computed over $N=100$ independent participants (one trial per participant). Model accuracy was computed over the same 100 stimuli per condition.
Because each participant contributed a single independent response and each stimulus was evaluated once per model, observations were treated as independent for the purpose of the proportion tests.
Two-proportion z-tests (pooled variance) were conducted in Python using the statsmodels library. All tests were two-tailed with $\alpha=0.05$.
C.2 VLM Evaluation Protocol
Vision–language models were evaluated using a single-call inference protocol. Most models were accessed via the OpenRouter API using an OpenAI-compatible interface, while Gemini models were accessed via their native provider interface.
Temperature was set to 0.0.
For each trial, video frames were sampled at a rate of 2 frames per second (FPS) across the full duration of the stimulus (up to a maximum of 10 frames per trial) and provided to the model as base64-encoded JPEG images.
Each model input consisted of the following components in fixed order:
1) Introductory task text explaining that the model would view frames from a short video. 2) The sampled video frames. 3) A prediction instruction asking the model to choose which of two candidate images would occur next. 4) The two candidate continuation images (A and B). 5) An explicit reminder of the required output format.
The prediction instruction explicitly required the model to respond using the keys prediction:, explanation:, and memory:, each on a separate line. The model was instructed to select either A or B as the prediction, provide a brief explanation of its reasoning, and name at least one object that appeared in the video.
Model outputs were parsed automatically using rule-based pattern matching to extract the predicted choice (A or B), the explanation text, and the memory response. The maximum generation length was set to 8192 tokens.